the pressure of a car tire is 310kPq.Id the pressure reading was made at 25°C, the reading of the tire pressure at -30°C would be?


1
Expert's answer
2021-03-22T06:32:25-0400

Assuming gas to be ideal,

"P_1=310KPa;T_1=25\degree C=273+25K=298K"

"P_2=?,T_2=-30\degree C=-30+273K=243K"

By using gay-Lussac's Law,"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

"\implies \frac{310}{298}=\frac{P_2}{243}"

"\implies P_2=252.78KPa"